# 3d Conformal Field Theories via Fuzzy Sphere Algebra

### v1

@@ Abstract
Fuzzy sphere models conjecturally realize 3d CFTs in small systems of spinful fermions, but why they work so well is still not fully understood. Their Hamiltonians are built from electron density operators projected to the lowest Landau Level. We analyze the algebra of the density modes and verify that it satisfies the Jacobi identity. The fuzzy sphere geometry admits two thermodynamic limits: a local planar limit yielding the fuzzy plane, and a commutative limit yielding an ordinary sphere. In the planar limit, high-angular-momentum modes recover the Girvin–MacDonald–Platzman algebra, whereas in the commutative limit the low-angular-momentum modes become semiclassical. We further find an explicit representation of the conformal algebra $so(3,2)$ in the minimal two-electron system and extend it to larger systems via an $so(3)$ equivariant coproduct. Because the coproduct splits one $so(3)$ representation into a tensor product, it is structurally mismatched with the thermodynamic limit of critical fuzzy sphere Hamiltonians.
@@ 1 Introduction
One 2+1d generalization of the anyonic chain is the fusion surface model Inamura and Ohmori (2024); Eck and Fendley (2025). Since we have not yet had success using them for approximating 3d CFTs, we instead propose to treat the fuzzy sphere Hamiltonians Zhu et al. (2023) as generalized spin chains (finite-dimensional quantum lattice systems in which we do not assume microscopic locality). The fuzzy sphere model describes spinful electrons on the 2-sphere with a monopole at the origin, projected to the lowest Landau level (LLL). Remarkably, it reproduces scaling dimensions of the 3d Ising CFT with good precision already in small systems of only $\mathcal{O}(10)$ electrons.
The Landau-level–projected density operators that enter the fuzzy-sphere Hamiltonians generate an intriguing algebra that is of independent interest, and which we study in this paper. After briefly reviewing the construction of fuzzy sphere Hamiltonians in Section 3, we connect the density modes both to fuzzy spherical harmonics from noncommutative field theory Madore (1992); Grosse et al. (1996) and to the matrix Lie algebra $su(2N)$ using their realization as fermionic bilinears (Section 4.1). Treating the commutator as defining an abstract algebra, we show that the bracket obeys the Jacobi identity and identify commuting subalgebras in 4.2. We then analyze two natural thermodynamic limits of the algebra: a local planar limit to the fuzzy plane and a commutative limit to an ordinary sphere in Sec. 5. The commutative, semiclassical regime is expected to be the more relevant one for the low-energy spectrum of the critical fuzzy-sphere Hamiltonians.
@@ 3.1 Hilbert space and fermionic density fields
$K(\Omega,\Omega^{\prime})=\frac{1}{R}\sum_{m=-s}^{s}Y^{(s)}_{s,m}(\Omega)\left(Y^{(s)}_{s,m}(\Omega^{\prime})\right)^{*}=\frac{2s+1}{R^{2}}e^{2is\alpha(\Omega,\Omega^{\prime})}\cos^{2s}{\left(\frac{\gamma}{2}\right)},\text{ and }$
@@ 3.2 Hamiltonian and symmetries
Beyond the Ising CFT, similar fuzzy sphere Hamiltonians have been used to realize a variety of 3d CFTs, including the free real scalar CFT He (2025); Taylor et al. (2025), the $O(N)$ and Wilson-Fisher CFTs Dey et al. (2025); Guo et al. (2025), DQCPs with emergent $SO(5)$ symmetry Zhou et al. (2024); Chen et al. (2024), the non-unitary Yang-Lee CFT Fan et al. (2025); Elias Miró and Delouche (2025); Arguello Cruz et al. (2025), a superconformal Ising CFT Tang et al. (2025), and the free fermion Zhou et al. (2025a) and Chern-Simons-matter CFTs Zhou and He (2025); Zhou et al. (2025b).
@@ 4.1 Commutation relations of the density modes
$\sum_{l,m}(-1)^{m}n^{\alpha}_{l,m}\,n^{\alpha}_{l,-m}=2n^{0}_{0,0}-\sum_{l,m}(-1)^{m}n^{0}_{l,m}\,n^{0}_{l,-m},\hskip 18.49988pt\alpha=x,y,z.$ (13)
@@ 4.2 Jacobi identity and commuting subalgebras
Does the mode algebra, viewed abstractly through the commutation relations (9), define a genuine Lie algebra? Let $\mathcal{A}_{s}=\{\tilde{n}^{A}_{l,m}\}_{l=0,|m|\leq l}$ be the algebra spanned by formal generators $\tilde{n}^{A}_{l,m}$, equipped with the bracket defined by (9). A priori this bracket need not satisfy Jacobi. To verify Jacobi, we use the fermionic realization. Let $\rho:\mathcal{A}_{s}\to\operatorname{End}(\mathcal{H}_{\rm Fock})$ be the map that sends the abstract generators to the corresponding lowest-Landau-level bilinears,
@@
Finally, a separate question is whether this algebra admits a nontrivial central extension; we briefly comment on this in Appendix C.
@@ 5 Thermodynamic limit of the density mode algebra
The fuzzy sphere geometry has two fundamentally different large-$s$ limits Chu et al. (2001); Steinacker (2024): On the one hand, if we zoom into a microscopic patch near the north pole, the algebra locally approaches the fuzzy plane: modes with angular momentum $l\sim\sqrt{s}$ have finite planar momentum $q\sim l/R$ and become generators of the GMP algebra. On the other hand, for small angular momenta $l\ll\sqrt{s}$, the non-commutativity vanishes as $s\to\infty$ and the fuzzy sphere becomes a commutative sphere. For the critical fuzzy-sphere Hamiltonians, we expect the low-energy sector to be controlled by this latter, semiclassical regime.
@@ 5.1 Quantum planar limit of large-$l$ density modes
$e^{i\mathbf{q}\cdot\hat{\mathbf{x}}}\,e^{i\mathbf{p}\cdot\hat{\mathbf{x}}}=\exp\!\Big(-\frac{i}{2}\Theta\,(\mathbf{q}\times\mathbf{p})\Big)\,e^{i(\mathbf{q}+\mathbf{p})\cdot\hat{\mathbf{x}}},\hskip 18.49988pt\mathbf{q}\times\mathbf{p}:=q_{1}p_{2}-q_{2}p_{1}.$ (22)
@@
$H\approx\rho^{x}_{\mathbf{0}}+\int d^{2}q\;V(q)\Big(\rho^{z}_{\mathbf{q}}\rho^{z}_{-\mathbf{q}}+\rho^{0}_{\mathbf{q}}\rho^{0}_{-\mathbf{q}}\Big),\hskip 18.49988ptV(q)=U_{0}+U_{1}q^{2}.$ (26)
@@ 5.2 Semiclassical limit of small-$l$ density modes
However, this naive semiclassical bracket does not directly reproduce several relations that appear to emerge in the low-energy sector of critical fuzzy-sphere Hamiltonians. For instance, Ref. He (2025) found that, upon restricting to states near the paramagnetic ground state, the operators $n^{+}_{l,m}$ and $n^{-}_{l,m}$ approximately behave as oscillator modes, with $n^{+}_{l,m},n^{-}_{l^{\prime},m^{\prime}}]\approx\delta_{l,l^{\prime}}\delta_{m+m^{\prime},0}n^{z}_{0,0}$. By contrast, the semiclassical commutator (28) generically yields a sum over many $L$ rather than collapsing to a single $L=0$ term. Similarly, the semiclassical commutator (27) does not by itself imply the conformal commutation relations observed numerically in the same low-energy subspace, e.g. $\Lambda_{z},[\Lambda_{z},D]]=4D$ in Fardelli et al. (2024), from the identifications $D\sim H$ and $\Lambda_{z}\sim\int d\Omega\,Y_{1,0}(\Omega)H(\Omega)$. Capturing these emergent relations likely requires a more refined scaling limit, as discussed in Sec. 2.4.
@@ 6.1 $so(3,2)$ algebra representation of $s=\tfrac{1}{2}$ density modes
In Appendix D, we show that this so(3,2)so(3,2)/$so(5)$ representation built from the $n^{A}_{l,m}$ density modes at $s=\tfrac{1}{2}$ does not extend to larger $s$, even if we allow an arbitrary enlargement of the internal spin degrees of freedom. For example, it is easy to see that the commutator $D_{+},D_{-}]$ produces an additional $n^{z}_{2,0}$ term when $s>\tfrac{1}{2}$.
@@ 6.2 Larger $so(3,2)$ representation from $so(3)$ equivariant coproduct
While this coproduct gives us a $so(3,2)$ representation defined on a larger Hilbert space, it is different from the thermodynamic limit of the critical fuzzy sphere models, which takes $s\to\infty$ within a single $V^{s}$. Using the standard coproduct of Lie algebras leads to a similar mismatch, see Appendix E. Nevertheless, it seems interesting that this coproduct induces a nontrivial $so(3,2)$ representation on a tensor product Hilbert space.
@@ Appendix B Derivation of the density mode commutator using fuzzy spherical harmonics
Appendix C Central extensions

### v2

@@ Abstract
Fuzzy sphere models conjecturally realize 3d CFTs in small systems of spinful fermions, but why they work so well is still not fully understood. Their Hamiltonians are built from electron density operators projected to the lowest Landau level. We analyze the algebra of the density modes and verify that it satisfies the Jacobi identity. The fuzzy sphere geometry admits two thermodynamic limits: a local planar limit yielding the fuzzy plane, and a commutative limit yielding an ordinary sphere. In the planar limit, high-angular-momentum modes recover the Girvin–MacDonald–Platzman algebra, whereas in the commutative limit, the low-angular-momentum modes become semiclassical. Upon further restricting to a subspace with few spin flips above the paramagnetic reference state, they behave approximately as harmonic oscillators. We also find an explicit representation of the conformal algebra $so(3,2)$ in the minimal two-electron system and extend it to larger systems via an $so(3)$ equivariant coproduct. Because the coproduct splits one $so(3)$ representation into a tensor product, it is structurally mismatched with the thermodynamic limit of critical fuzzy sphere models.
@@ 1 Introduction
One 2+1d generalization of the anyonic chain is the fusion surface model Inamura and Ohmori (2024); Eck and Fendley (2025). Since we have not yet had success using them for approximating 3d CFTs, we instead propose to treat the fuzzy sphere Hamiltonians Zhu et al. (2023) as generalized spin chains (finite-dimensional quantum lattice systems in which we do not assume microscopic locality). The fuzzy sphere model describes spinful electrons on the 2-sphere with a monopole at the origin, projected to the lowest Landau level (LLL). Within a two-parameter family of $SO(3)$ invariant Hamiltonians, one can tune to a critical point whose low-lying spectrum matches 3d Ising CFT scaling dimensions with good precision already for $\mathcal{O}(10)$ electrons. We emphasize, however, that obtaining CFT-like spectra from small 2+1d systems is not unique to the fuzzy sphere: similar behavior occurs in more conventional discretizations, such as the 2+1d transverse-field Ising model on the icosahedron Lao and Rychkov (2023), where good accuracy requires conformal perturbation theory. Conformal perturbation theory is likewise useful on the fuzzy sphere to further improve the extracted scaling dimensions Läuchli et al. (2025); Voinea et al. (2025). A central conceptual problem is locating the conformal points of the Hamiltonian. Possible routes include conformal perturbation theory and enlarged symmetry, but we leave this challenging question to the future.
The Landau-level–projected density operators that enter the fuzzy-sphere Hamiltonians generate an intriguing algebra that is of independent interest, and which we study in this paper. After briefly reviewing the construction of fuzzy sphere Hamiltonians in Section 3, we connect the density modes both to fuzzy spherical harmonics from noncommutative field theory Madore (1992); Grosse et al. (1996) and to the matrix Lie algebra $su(2N)$ using their realization as fermionic bilinears (Section 4.1). Treating the commutator as defining an abstract algebra, we show that the bracket obeys the Jacobi identity and identify commuting subalgebras in 4.2. We then analyze two natural thermodynamic limits of the algebra: a local planar limit to the fuzzy plane and a commutative limit to an ordinary sphere in Sec. 5. The commutative, semiclassical regime is expected to be the more relevant one for the low-energy spectrum of the critical fuzzy-sphere Hamiltonians. We also refine the observation of He (2025) that, after restricting to a low-excitation subspace above the paramagnetic state, the density modes admit an approximate harmonic-oscillator description.
@@ 3.1 Hilbert space and fermionic density fields
$K(\Omega,\Omega^{\prime})=\frac{1}{R^{2}}\sum_{m=-s}^{s}Y^{(s)}_{s,m}(\Omega)\left(Y^{(s)}_{s,m}(\Omega^{\prime})\right)^{*}=\frac{2s+1}{R^{2}}e^{2is\alpha(\Omega,\Omega^{\prime})}\cos^{2s}{\left(\frac{\gamma}{2}\right)},\text{ and }$
@@ 3.2 Hamiltonian and symmetries
Beyond the Ising CFT, similar fuzzy sphere Hamiltonians have been used to realize a variety of 3d CFTs, including the free real scalar CFT He (2025); Taylor et al. (2025), the $O(N)$ and Wilson-Fisher CFTs Dey et al. (2025); Guo et al. (2025), DQCPs with emergent $SO(5)$ symmetry Zhou et al. (2024); Chen et al. (2024), the non-unitary Yang-Lee CFT Fan et al. (2025); Elias Miró and Delouche (2025); Arguello Cruz et al. (2025), a superconformal Ising CFT Tang et al. (2025), and the free fermion Zhou et al. (2025a); Voinea et al. (2026) and Chern-Simons-matter CFTs Zhou and He (2025); Zhou et al. (2025b). In addition, fractional-filling fuzzy-sphere models (rather than integer $\nu=1$ filling) can also realize 3d CFTs Voinea et al. (2025).
@@ 4.1 Commutation relations of the density modes
$\sum_{l,m}(-1)^{m}n^{\alpha}_{l,m}\,n^{\alpha}_{l,-m}=2(2s+1)n^{0}_{0,0}-\sum_{l,m}(-1)^{m}n^{0}_{l,m}\,n^{0}_{l,-m},\hskip 18.49988pt\alpha=x,y,z.$ (13)
@@ 4.2 Jacobi identity and commuting subalgebras
Does the mode algebra, viewed abstractly through the commutation relations (9), define a genuine Lie algebra? Let $\mathcal{A}_{s}=\{\tilde{n}^{A}_{l,m}\}_{0\leq l=0\leq 2s,|m|\leq l}$ be the algebra spanned by formal generators $\tilde{n}^{A}_{l,m}$, equipped with the bracket defined by (9). A priori this bracket need not satisfy Jacobi. To verify Jacobi, we use the fermionic realization. Let $\rho:\mathcal{A}_{s}\to\operatorname{End}(\mathcal{H}_{\rm Fock})$ be the map that sends the abstract generators to the corresponding lowest-Landau-level bilinears,
@@
Finally, a separate question is whether this algebra admits a nontrivial central extension; we briefly comment on this in Appendix D.
@@ 5 Thermodynamic limit of the density mode algebra
The fuzzy sphere geometry has two fundamentally different large-$s$ limits Chu et al. (2001); Steinacker (2024): On the one hand, if we zoom into a microscopic patch near the north pole, the algebra locally approaches the fuzzy plane: modes with angular momentum $l\sim\sqrt{s}$ have finite planar momentum $q\sim l/R$ and become generators of the GMP algebra. On the other hand, for small angular momenta $l\ll\sqrt{s}$, the non-commutativity vanishes as $s\to\infty$ and the fuzzy sphere becomes a commutative sphere. For the critical fuzzy-sphere Hamiltonians, we expect the low-energy sector to be controlled by this latter, semiclassical regime. Moreover, Sec. 5.3 shows that if one additionally focuses on a low-excitation sector consisting of a few low-momentum spin flips above the paramagnetic reference state, the density modes admit an approximate harmonic-oscillator description.
@@ 5.1 Quantum planar limit of large-$l$ density modes
$e^{i\mathbf{q}\cdot\hat{\mathbf{x}}}\,e^{i\mathbf{p}\cdot\hat{\mathbf{x}}}=\exp\!\Big(-\frac{i}{2}\Theta\,(\mathbf{q}\times\mathbf{p})\Big)\,e^{i(\mathbf{q}+\mathbf{p})\cdot\hat{\mathbf{x}}},\hskip 18.49988pt\mathbf{q}\times\mathbf{p}:=q_{1}p_{2}-q_{2}p_{1}.$
@@
$H\approx-h\rho^{x}_{\mathbf{0}}+\int d^{2}q\;V(q)\Big(\rho^{0}_{\mathbf{q}}\rho^{0}_{-\mathbf{q}}-\rho^{z}_{\mathbf{q}}\rho^{z}_{-\mathbf{q}}\Big),\hskip 18.49988ptV(q)=U_{0}+U_{1}q^{2}.$ (25)
@@ 5.2 Semiclassical limit of small-$l$ density modes
However, this naive semiclassical bracket does not directly reproduce several relations that appear to emerge in the low-energy sector of critical fuzzy-sphere Hamiltonians. For example, the semiclassical commutator (26) does not by itself imply the conformal commutation relations observed numerically in the same low-energy subspace, e.g. $\Lambda_{z},[\Lambda_{z},D]]=4D$ in Fardelli et al. (2024), from the identifications $D\sim H$ and $\Lambda_{z}\sim\int d\Omega\,Y_{1,0}(\Omega)H(\Omega)$. Capturing these emergent relations likely requires a more refined scaling limit, as discussed in Sec. 2.4.
5.3 Oscillator approximation in subspace above paramagnetic state
[…]
The commutators $n^{+}_{l,m},n^{+}_{l^{\prime},m^{\prime}}]=[n^{-}_{l,m},n^{-}_{l^{\prime},m^{\prime}}]=0$ hold exactly, while the mixed commutator is given by
$\left[n^{+}_{l,m},n^{-}_{l^{\prime},m^{\prime}}\right]=$ $\sum_{\begin{subarray}{c}L=0\\ l+l^{\prime}+L\text{ even}\end{subarray}}h_{L}(s,l,l^{\prime})(-1)^{m+m^{\prime}}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}n^{z}_{L,m+m^{\prime}}$ (28)
- $\sum_{\begin{subarray}{c}L=0\\ l+l^{\prime}+L\text{ odd}\end{subarray}}h_{L}(s,l,l^{\prime})(-1)^{m+m^{\prime}}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}n^{0}_{L,m+m^{\prime}}.$
Following He (2025), we consider the subspace spanned by states with a small number $k$ of spin flips and bounded angular momentum,
[…]
On this subspace, the density mode $n^{z}_{0,0}$ is extensive, with eigenvalue N+2k-N+2k on a state with $k$ spin flips. In contrast, for $l>0$ the density modes satisfy the operator norm bound (see Appendix C for derivation)
$\|n^{z}_{l,m}\|_{P_{k}}\leq 2k\sqrt{2l+1},$
where the $P_{k}$ subscript denotes the restriction to the sector with $k$ up spins. The same bound holds for $n^{0}_{l,m}$.
Although (28) contains many $n^{z}_{L,m}$ terms, on the subspace of $k$ spin flips with $l_{i}\leq l_{\text{max}}$ only those with $L\leq 2kl_{\text{max}}$ survive. The reason is that $n^{z}_{L,m}$ carries angular momentum $L$, so for $L>2kl_{\text{max}}$ it necessarily takes any state inside the subspace (29) to states with angular momentum greater than $kl_{\text{max}}$ outside the subspace. Therefore, the operator norm of the remaining contribution can be bounded as
$\left\|\sum_{L=1}^{2kl_{\text{max}}}h_{L}(s,l,l^{\prime})(-1)^{m+m^{\prime}}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}n^{z}_{L,m+m^{\prime}}\right\|_{P_{k}}$
[…]
for some constants $C$, $c$ independent of $s$. The inequality is derived in Appendix C.
In particular, if $l,l^{\prime}\leq l_{\text{max}}$ and $kl_{\text{max}}=o(\sqrt{s})$, then the right-hand side divided by $N$ vanishes as $s\to\infty$. Then the rescaled modes satisfy the oscillator approximation on this subspace,
$\Bigg[\frac{n^{+}_{l,m}}{\sqrt{N}},\,\frac{n^{-}_{l^{\prime},m^{\prime}}}{\sqrt{N}}\Bigg]\;\approx\;\delta_{l,l^{\prime}}\,\delta_{m+m^{\prime},0}\,\frac{n^{z}_{0,0}}{N}\;\approx\;\delta_{l,l^{\prime}}\,\delta_{m+m^{\prime},0}.$ (30)
Sharpening the bounds on the $3j$ and $6j$ symbols used in the derivation in Appendix C would likely improve the dependence on $k$ and $l_{\text{max}}$, but we do not pursue this here.
@@ 6.1 $so(3,2)$ algebra representation of $s=\tfrac{1}{2}$ density modes
In Appendix E, we show that this so(3,2)so(3,2)/$so(5)$ representation built from the $n^{A}_{l,m}$ density modes at $s=\tfrac{1}{2}$ does not extend to larger $s$, even if we allow an arbitrary enlargement of the internal spin degrees of freedom. For example, it is easy to see that the commutator $D_{+},D_{-}]$ produces an additional $n^{z}_{2,0}$ term when $s>\tfrac{1}{2}$.
@@ 6.2 Larger $so(3,2)$ representation from $so(3)$ equivariant coproduct
While this coproduct gives us a $so(3,2)$ representation defined on a larger Hilbert space, it is different from the thermodynamic limit of the critical fuzzy sphere models, which takes $s\to\infty$ within a single $V^{s}$. Using the standard coproduct of Lie algebras leads to a similar mismatch, see Appendix F. Nevertheless, it seems interesting that this coproduct induces a nontrivial $so(3,2)$ representation on a tensor product Hilbert space.
@@ Appendix B Derivation of the density mode commutator using fuzzy spherical harmonics
Appendix C Derivation of oscillator approximation
[…]
In the $k=1$ subspace spanned by states $|\ket{a;b}:=c^{\dagger}_{a,\uparrow}c_{b,\downarrow}\ket{\Downarrow}$, the matrix elements of $n^{z}_{l,m}$ with $l>0$ are
[…]
$=\mathcal{N}_{s,l}\left((\hat{Y}_{l,m})_{a^{\prime},a}\delta_{b,b^{\prime}}+(\hat{Y}_{l,m})_{b,b^{\prime}}\delta_{a,a^{\prime}}\right)$
Then using the triangle inequality for the two terms, we can bound the operator norm of $n^{z}_{l,m}$ with $l>0$ in this subspace by
$\|n^{z}_{l,m}\|\leq 2\mathcal{N}_{s,l}\|\hat{Y}_{l,m}\|.$
To bound the operator norm of $\hat{Y}_{l,m}$, we use that the norm is equal to the square root of the largest eigenvalue of $\hat{Y}^{\dagger}_{l,m}\hat{Y}_{l,m}$. This is a diagonal matrix,
$(\hat{Y}^{\dagger}_{l,m}\hat{Y}_{l,m})_{bb^{\prime}}=\sum_{a}(\hat{Y}_{l,m})^{*}_{ab}(\hat{Y}_{l,m})_{ab^{\prime}}=\delta_{b,b^{\prime}}\sum_{a}|(\hat{Y}_{l,m})_{ab}|^{2}=\delta_{b,b^{\prime}}|(\hat{Y}_{l,m})_{b+m,b}|^{2},$
because $(\hat{Y}_{l,m})_{ab}$ is nonzero only when b=mab=m-a and $(\hat{Y}_{l,m})_{ab^{\prime}}$ is nonzero only when $b^{\prime}=m-a$. Therefore, the operator norm of $\hat{Y}_{l,m}$ is equal to
$\|\hat{Y}_{l,m}\|=\max_{b}|(\hat{Y}_{l,m})_{b+m,b}|=\sqrt{2l+1}\max_{b}\left|\begin{pmatrix}s&s&l\\ b+m&-b&-m\end{pmatrix}\right|\leq\sqrt{\frac{2l+1}{2s+1}}.$
The last inequality follows from the condition
$(2j_{3}+1)\sum_{m_{1},m_{2}}\begin{pmatrix}j_{1}&j_{2}&j_{3}\\ m_{1}&m_{2}&m_{3}\end{pmatrix}^{2}=\begin{cases}1&|j_{1}-j_{2}|\leq j_{3}\leq j_{1}+j_{2}\\ 0&\text{else }\end{cases},$
which implies
$\begin{pmatrix}j_{1}&j_{2}&j_{3}\\ m_{1}&m_{2}&m_{3}\end{pmatrix}^{2}\leq\frac{1}{2\max{\{j_{1},j_{2},j_{3}\}}+1}.$
Together with the fact that $\mathcal{N}_{s,l}=(2s+1)\begin{pmatrix}s&s&l\\ s&-s&0\end{pmatrix}\leq\sqrt{2s+1}$, this leads to
$\|n^{z}_{l,m}\|\leq 2\sqrt{2l+1}.$
In contrast, for $n^{z}_{0,0}$ we have
[…]
therefore $n^{z}_{0,0}$ is extensive in $N$ while $\|n^{z}_{l,m}\|=\mathcal{O}(\sqrt{l})$ for $l>0$ (and analogously for $n^{0}_{l,m}$).
In the sector with $k$ spin flips spanned by states $|\ket{A;B}:=c^{\dagger}_{a_{1}}\dots c^{\dagger}_{a_{k}}c_{b_{1}}\dots c_{b_{k}}\ket{\Downarrow}$ with $A=\{a_{1},\dots,a_{k}\}$ and $B=\{b_{1},\dots,b_{k}\}$, the matrix elements of $n^{z}_{l,m}$ are
[…]
Using again the triangular inequality for the $2k$ terms, the operator norm is then bounded as
$\|n^{z}_{l,m}\|\leq 2k\mathcal{N}_{s,l}\|\hat{Y}_{l,m}\|\leq 2k\sqrt{2l+1}.$
Again, the same holds for $n^{0}_{l,m}$. In contrast, the operator norm of $n^{z}_{0,0}$ is again extensive, concretely
[…]
$\left\|\sum_{L=1}^{2kl_{\text{max}}}h_{L}(s,l,l^{\prime})(-1)^{m+m^{\prime}}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}n^{z}_{L,m+m^{\prime}}\right\|_{P_{k}}$
$\leq\left(\sum_{L=1}^{2kl_{\text{max}}}(h_{L}(s,l,l^{\prime}))^{2}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}^{2}\right)^{1/2}\left(\sum_{L=1}^{2kl_{\text{max}}}\|n^{z}_{L,m+m^{\prime}}\|^{2}_{P_{k}}\right)^{1/2}$
To bound the first factor, we use $\begin{pmatrix}j_{1}&j_{2}&j_{3}\\ m_{1}&m_{2}&m_{3}\end{pmatrix}\leq\frac{1}{\sqrt{2\max{\{j_{1},j_{2},j_{3}\}}+1}}$ on three different 3j symbols,
$\sum_{L=1}^{2kl_{\text{max}}}(h_{L}(s,l,l^{\prime}))^{2}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}^{2}\leq(2l+1)(2l^{\prime}+1)\sum_{L=1}^{2kl_{\text{max}}}\frac{\begin{Bmatrix}l&l^{\prime}&L\\ s&s&s\end{Bmatrix}^{2}}{\begin{pmatrix}s&s&L\\ s&-s&0\end{pmatrix}^{2}}.$
Orthogonality of the 6j symbols then gives
$\sum_{L=1}^{2kl_{\text{max}}}\frac{\begin{Bmatrix}l&l^{\prime}&L\\ s&s&s\end{Bmatrix}^{2}}{\begin{pmatrix}s&s&L\\ s&-s&0\end{pmatrix}^{2}}\leq\frac{1}{2s+1}\max_{L=1,\dots,2kl_{\text{max}}}\begin{pmatrix}s&s&L\\ s&-s&0\end{pmatrix}^{-2}$
The inverse squared 3j symbol is known to be
$\begin{pmatrix}s&s&L\\ s&-s&0\end{pmatrix}^{-2}=\frac{(2s-L)!(2s+L+1)!}{((2s)!)^{2}}.$
For $L_{\text{max}}=2kl_{\text{max}}\leq s/2$, this can be bounded using Stirling’s approximation as
[…]
where $c$ and $C$ are some constants. If $L_{\text{max}}\ll\sqrt{s}$, this exponential is $\mathcal{O}(1)$. Altogether, this gives the bound
$\left\|\sum_{L=1}^{2kl_{\text{max}}}h_{L}(s,l,l^{\prime})(-1)^{m+m^{\prime}}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}n^{z}_{L,m+m^{\prime}}\right\|_{P_{k}}$
[…]
Appendix D Central extensions